Find the x intercept of these parabola by completing the square
a)y=x^2-4x+1
b)y=2x^2-12x-7
Can you please help me out? Thanks so much in advance:)
Can you please show all the steps it would really help me understand better thank you:)

Answer by waynest(281)

x^2 -4x + 1
x^2 -4x + (-4/2)^2 - (-4/2)^2 + 1
x^2 -4x + 4 - 4 + 1
(x - 2)^2 - 3
(x - 2)^2 = 3
sqrt(x - 2)^2 = +-sqrt3
x - 2 = +- sqrt3
x = 2+- sqrt3
x = 2+ sqrt 3
x = 3.73
x = 2 - sqrt3
x = 0.27
2x^2 -12x -7
divide by 2
x^2 - 6x -3.5
x^2 - 6x + (-6/2)^2 - (-6/2)^2 - 3.5
x^2 - 6x + 9 -9 -3.5
(x - 3)^2 -12.5 = 0
(x - 3)^2 = 12.5
sqrt(x - 3)^2 = +-sqrt12.5
x - 3 = +- sqrt12.5
x = 3 +- sqrt12.5
x = 3 + sqrt12.5
x = 6.54
x = 3 - sqrt12.5
x = -0.54